OpenStudy (anonymous):

What is the inverse of ƒ? ƒ(x) = 8x

OpenStudy (anonymous):

put y = f(x)

OpenStudy (anonymous):

y = 8x x = y/8 so f^-1 (x) = x/8

OpenStudy (anonymous):

THANK YOU SO MUCH !!!!

OpenStudy (anonymous):

Yes he is right \[f ^{-1}(x)=\frac{ x }{ }\] is the inverse

OpenStudy (anonymous):

oops x/8*